If a=phi, b={1,2}, c={-1,-2}, then a×b×c has 4 elements.
Is the statement true or false? Justify your answer.
1
Expert's answer
2019-05-14T09:49:01-0400

By the definition,

In set theory (and, usually, in other parts of mathematics), a Cartesian product is a mathematical operation that returns a set (or product set or simply product) from multiple sets. That is, for sets A and B, the Cartesian product A × B is the set of all ordered pairs (ab) where a ∈ A and b ∈ B. Products can be specified using set-builder notation, e.g.


"A\times B=\left\{(a,b)\left|a\in A\quad and \quad b\in B\right.\right\}"

By the definition,

The cardinality of a set is the number of elements of the set. The cardinality of the output set is equal to the product of the cardinalities of all the input sets. That is,


"\left|A\times B\right|=\left|A\right|\cdot\left|B\right|"

Similarly,


"\left|A\times B\times C\right|=\left|A\right|\cdot\left|B\right|\cdot\left|C\right|"

and so on.

( More information: https://en.wikipedia.org/wiki/Cartesian_product )

In our case,


"\left\{\begin{array}{l}
A=\left\{\phi\right\}\rightarrow\left|A\right|=1\\
B=\left\{1,2\right\}\rightarrow\left|B\right|=2\\
C=\left\{-1,-2\right\}\rightarrow\left|C\right|=2\\
\end{array}\right."

Then,


"\left|A\times B\times C\right|=\left|A\right|\cdot\left|B\right|\cdot\left|C\right|=1\cdot 2\cdot 2=4"

Conclusion,


"\boxed{A\times B\times C\,\,\,has\,\,\, 4\,\,\, elements.\,\,\,The\,\,\,statement\,\,\, is\,\,\, TRUE}"